# Multilayer networks characterize human-mobility patterns by industry sector for the 2021 Texas winter storm

### v1

@@ 2 Methods
2.1 Construction of a network that encodes movements between census tracts (CTs).
@@
We study a dataset of weekly human movements from the data provider SafeGraph. The approximately 700 GB dataset is collected based on the GPS locations of opt-in smart mobile devices (mostly smartphones), and captures weekly movements of people from a home location, recorded at the corresponding Census Block Group (CBG), to a destination location, i.e., a specific Point of Interest (POI). POIs are identified using SafeGraph’s Placekey system, a universal location identifier that combines a geospatial encoding system with a unique POI identifier that provides information including the name, latitude and longitude, business details, and industry classification. For privacy reasons, SafeGraph omits sparse data in which fewer than four visitors are recorded in a given week from any home CBG to a POI.
Our study area is Harris County, TX which was severely affected by the 2021 winter storm. Figure 1(a) shows a visualization of example SafeGraph data encoding the movements of people to hospitals in Harris County. Observe that the original SafeGraph data can be encoded by a time-varying, bipartite graph $G(t)$, for $t=1,\dots,T$, where $T$ is the total number of weeks studied. Each graph $G(t)$ is composed of weighted, directed edges that encode the number of observed movements between a source CBG and destination POI. We denote $\mathcal{H}=\{CBG_{i}\}_{i=1}^{H}$ to be a set of source nodes (i.e., the home CBGs associated with mobile devices), where $H$ is the total number of CBGs in the studied area, and let $\mathcal{P}=\{POI_{i}\}_{i=1}^{P}$ be a set of destination nodes (i.e., the set of POIs), where $P$ is the total number of POIs. Each edge $(i,j)$ in $G(t)$ has a weight $B_{ij}(t)$ that encodes the number of observed movements from $CBG_{i}$ to $POI_{j}$ during week $t$. Equivalently, each graph $G(t)$ can be encoded by a weighted adjacency matrix, $B(t)\in\mathbb{R}^{H\times P}$. We use the Python library safegraph_py to process and prepare the data into this graph-structured format and computations were implemented on the NCAR-Wyoming Supercomputing Center.
Figure 1: Example SafeGraph movement data and spatial aggregation across census tracts (CTs). (a) Visualization of observed movements from home neighborhoods to hospitals in Harris County, TX for the week ending February 22, 2021 (i.e., the storm week). Home locations are recorded using U.S. Census Block Groups (CBGs) $\mathcal{H}=\{CBG_{i}\}_{i=1}^{H}$, whereas destinations locations are Points of Interests $\mathcal{P}=\{POI_{i}\}_{i=1}^{P}$ with known latitudes, longitudes, and other information such as industry category (e.g. hospitals). The movements are naturally represented by a weighted, bipartite network with a time-varying adjacency matrix $B(t)\in\mathbb{R}^{H\times P}$. (b) The constructed network encoded by a time-varying adjacency matrix in which $A_{ij}(t)$ encodes movements from home CBGs in $CT_{i}$ to POIs in $CT_{j}$ during week $t$.
@@
Following the literature [34, 35, 29], we aggregate CBGs to census tracts (CT) and construct the network based on CTs and POIs. See Figure 1(b) for illustration. More precisely, we define $\mathcal{C}=\{CT_{i}\}_{i=1}^{C}$ to be a set of CTs of interest, where $C$ is the total number of CTs ($C=786$ for Harris County) and let $\mathcal{H}_{i}$ and $\mathcal{P}_{i}$ denote, respectively, the CBGs and POIs within $CT_{i}$. Then the combined observed movement from CBGs in $CT_{i}$ to POIs in $CT_{j}$ during week $t$ is given by
@@
The remainder of this study examines time-varying networks encoded by square adjacency matrices $A(t)$ that are size $C\times C$. Finally, for each network, we define $d^{out}_{i}(t)=\sum_{j}A_{ij}(t)$ to be the out-degree—that is, a measure for all movements during week $t$ that leave $CT_{i}$—and $d^{in}_{j}(t)=\sum_{i}A_{ij}(t)$ to be the in-degree—that is, a measure for all movements to POIs within $CT_{j}$ during week $t$. For the mobility network in Figure 1(b), for example, in- and out-degrees for $CT_{1}$ would be $d^{in}_{1}(t)=A_{11}(t)$ and $d^{out}_{1}(t)=A_{11}(t)+A_{12}(t)+A_{13}(t)$.
2.2 Stratification of the constructed network into multiple layers encoding behavioral categories
Herein, we refer to the act of partitioning a network’s edges into categorized sets of edges associated with network layers as “stratification” [36]. We construct behavior-stratified multilayer networks in which different network layers encode different behavioral categories of movement (e.g., visits to schools, hospitals, etc.). More specifically, we study a type of multilayer network called a multiplex network in which each layer consists of the same set of nodes (i.e., in our case, the set of CTs). We classify behavioral categories of movement based on the 2017 North American Industry Classification System (NAICS), which were used to classify POIs in the SafeGraph data.
Importantly, NAICS is a hierarchical categorization scheme, allowing us to stratify movement data into a hierarchical set of mobility network layers. See Figure 2(a) for an toy illustration of this hierarchy of network layers. That is, at a coarse level of the hierarchy, the network encoding all movements can be stratified into into layers encoding movements to Educational Services (NAICS code 61), Health Care and Social Services (62), Accommodation and Food Service (72), and additional categories. Whereas at a finer level of the hierarchy, the network layer encoding movements to Health Care and Social Services can be further stratified into network “sublayers” that encode, e.g., movements to locations associated with Ambulatory Health Care Facilities (621), Hospitals (622), and so on. In Figure 2(b), we depict a map of CTs in Harris County TX overlaid with visualizations of example network layers at 3 different levels of coarseness for the movement categories: all movements (top), movements to Health Care and Social Assistance location (middle), and movements to Hospitals (lower). In our study, the three-digit layer provides a sufficient breakdown for fine stratification, except for educational services where we consider four digits of NAICS codes.
Figure 2: POI classifications stratify a mobility network into a hierarchy of network layers. (a) Toy illustration for the hierarchical stratification of a mobility network into network layers that encode different behavioral categories of movements. The number of digits in a NAICS code determines the hierarchy depth (i.e., level of coarseness when refining movements categories into subcategories). (b) A map of CTs in Harris County (bottom) overlaid by three example networks at three different coarseness levels: all movements (top), health care and social assistance (middle), and hospitals (lower).
@@ 2.2 Stratification of the constructed network into multiple layers encoding behavioral categories
Each NAICS category has a numerical code with 2 to 6 digits depending on level in the hierarchy, with two digits at the coarsest level and six digits at the finest, most-granular level. At the coarsest level of the hierarchy, each NAICS category has a 2-digit numerical code indicating whether the POI is associated with one of twenty-four industry classifications: {62: Health Care and Social Assistance, 72: Accommodation and Food Services, 61: Educational Services, …}. Letting $\mathcal{N}$ denote the set of NAICS codes with a fixed number of digits, for each $n\in\mathcal{N}$ we define $\mathcal{P}^{(n)}_{j}$ as the set of POIs within $CT_{j}$ having NAICS code $n$. The network layers’ adjacency matrices are then obtained by $A^{(n)}_{ij}(t)=\sum_{i^{\prime}\in\mathcal{H}_{i},j^{\prime}\in\mathcal{P}^{(n)}_{j}}B_{i^{\prime}j^{\prime}}(t).$ For each network layer $n$ and $CT_{i}$, we define the time-varying in- and out-degrees by $d^{in,(n)}_{i}(t)=\sum_{j}A^{(n)}_{ji}(t)$ and $d^{out,(n)}_{i}(t)=\sum_{j}A^{(n)}_{ij}(t)$, respectively. Finally, note that summing $A^{(n)}_{ij}(t)$ over all possible $n\in\mathcal{N}$ recovers the adjacency matrix $\sum_{n}A^{(n)}(t)=A(t)$ of the original, non-stratified network (i.e., also called the layer-aggregated network). One can similarly obtain the in- and out-degrees for the network encoding all movements by summing over the network layers: $d^{in}_{i}(t)=\sum_{n}d^{in,(n)}_{i}(t)$ and $d^{out}_{i}(t)=\sum_{n}d^{out,(n)}_{i}(t)$.
Figure 3: Breakdown of POIs and observed movements into NAICS categories. (a) Fraction of POIs in each NAICS category for Harris County. (b) Fraction of total observed movements in each NAICS category. In both panels, the outer rings show the categorical stratification according to a coarse scale with 2-digit NAICS codes, whereas the inner rings show a finer stratification using NAICS codes either three or four digits. See Appendix 5.1 for a detailed description for the stratification of categories into subcategories.
@@ 2.3 Z-scores quantify movement-change significance
We quantify the storm’s impact on total movements for different network layers by comparing the total movements $m^{(n)}(t)$ during the storm week to a baseline of weekly movement, and we use z-scores to measure deviation from typical behavior. We apply this approach to different network layers to identify which movement categories undergo statistically significant change. After visually inspecting time series encoding weekly movements for several years of data (see Appendix 5.2), we select the baseline that consists of the six weeks prior to the storm, $\mathcal{T}_{base}=\{2,3,\dots,7\}$. Letting $\mu$ and $\sigma$ denote the mean and standard deviation of $m(t)$ across $t\in\mathcal{T}_{base}$, we compute the z-score $Z(t)=\frac{m(t)-\mu}{\sigma}$ to study the storm’s impact on all movement categories. Similarly, for each movement category $n$ we let $\mu^{(n)}$ and $\sigma^{(n)}$ denote the mean and standard deviation of $m^{(n)}(t)$ across $t\in\mathcal{T}_{base}$ and compute the z-score
@@
A visualization of this calculation is provided in Figure 4 for the network layer encoding movements to locations associated with Healthcare and Social Assistance (i.e., NAICS code 62). We find $Z^{(62)}(8)\approx-27$, implying that these movements underwent a significant decrease during the storm, i.e., by roughly 27 standard deviations. Note that week 8 is the storm week. We compare z-scores across movement categories in Section 3.1.
Figure 4: Z-scores quantify movement change. (a) For the network layer encoding movements to locations associated with Healthcare and Social Assistance (i.e., NAICS code 62), we show the total movements $m^{(62)}(t)$ versus $t$. Red and gray shading highlight the storm week and the weeks $\mathcal{T}_{base}$ used to construct a baseline. (b) To quantify the change in movements during the storm week, we calculate the z-score $Z^{(62)}(8)\approx-27$ using a baseline mean, $\mu^{(62)}$, and standard deviation, $\sigma^{(62)}$.
@@ 2.4 Regression analysis relates movements to infrastructure, demographic, and socioeconomic information
In Section 3.3, we study the in- and out-degrees for network layers and investigate how these structural properties relate to infrastructure information (i.e., derived from the POIs) as well social factors including demographic and socioeconomic information. To this end, we gathered data from the U.S. Census Bureau, accessing tables from the 2010 American Community Survey and filtering for year 2019 and CTs in Harris County, TX. For each CT, we assembled data for 13 social factors: population (B01003), population density, under 18 (DP05), under 5 (DP05), income (B19013), unemployment (DP03), poverty rate (S0601), non-white percentage (B02001), non-hispanic and non-black percentage (S0601), owner occupied percentage (B25003), renter occupied percentage (B25003), education level (S0601). Four key social factors are visualized in Figure 5.
Figure 5: Socioeconomic and demographic data from U.S. Census. Choropleth maps display demographic and socioeconomic factors across CT in Harris County, TX obtained from the U.S. Census. The maps show the distribution of population, income, percentage of non-white population, and poverty rate, offering a spatial overview of these key characteristics. CT 980000 is colored white, because the population is so small (i.e., 4 people).
To identify which social factors have the strongest correlation with the networks’ in- and out-degrees, we conduct a multilinear regression. Noting that some social factors are correlated, provide redundant information, and cause regression instability, we sought to obtain a smaller set of social factors. Specifically, we conducted a variance inflation factor (VIF) test to identify and remove variables having multicollinearity, which can distort regression coefficients and reduce the model’s reliability. For each social factor, $X^{(i)}\in\mathbb{R}^{786},i=1,\dots,13$, we calculated $VIF_{i}=1/(1-R^{2}_{i})$, where
@@
$R^{2}_{i}$ is called the coefficient of determination. Here, each $\hat{X}^{(i)}$ is a predicted value from regression and $\bar{X}^{(i)}$ is the mean across observed values. We successively removed variables with the highest VIF (implemented using the statsmodels module in Python) until all VIFs were less than five, leaving the following six social factors and VIFs shown in Table 1.
$X^{(i)}$ Description $VIF_{i}$
$X^{(1)}$ Population 3.38
$X^{(2)}$ Population Density 2.67
$X^{(3)}$ Income 2.67
$X^{(4)}$ Non-White Percentage 4.85
$X^{(5)}$ Poverty Rate 3.93
$X^{(6)}$ Unemployment Rate 4.25
Table 1: Variance inflation factors (VIFs) for 6 selected socioeconomic and demographic variables.
In addition to the social features above, we also include infrastructure features for each CT in the form of number of POIs. We look at total number of POIs for each CT as well as a breakdown of number of POIs for the high-movement categories, $\mathcal{N}_{high}$, shown in Figure 3. We highlight that we never include both the total number of POIs and the stratified POIs into categories, since that would introduce collinearity into the model (i.e., the total of POIs equals the summation over POIs in different categories). Each model defines a relationship between either in- or out-degrees and a set of features, $\left\{X^{(i)}\mid i\in S\right\}$, where $S$ is a set of select indexed features (possibly including demographic, socioeconomic, and infrastructure information). We then fit linear regression models of the form

### v2

@@ 4 Methods
4.1 Construction of multilayer networks
We study a dataset of weekly human movements from the data provider SafeGraph. The approximately 700 GB dataset is collected based on the GPS locations of opt-in smart mobile devices (mostly smartphones), and captures weekly movements of people from a home location, recorded at the corresponding census block group, to a destination location, i.e., a specific point of interest (POI). We use the Python library safegraph_py to process and prepare the data into this graph-structured format and computations were implemented on the NCAR-Wyoming Supercomputing Center.
@@ 4.1 Construction of multilayer networks
The original SafeGraph data can be encoded by a time-varying, bipartite graph $G(t)$, for $t=1,\dots,T$, where $T$ is the total number of weeks studied. Each graph $G(t)$ is composed of weighted, directed edges that encode the number of observed movements between a source census block group and destination POI. We denote $\mathcal{H}=\{CBG_{i}\}_{i=1}^{H}$ to be a set of source nodes (i.e., the home census block groups associated with mobile devices), where $H$ is the total number of census block groups in the studied area, and let $\mathcal{P}=\{POI_{i}\}_{i=1}^{P}$ be a set of destination nodes (i.e., the set of POIs), where $P$ is the total number of POIs. POIs are identified using SafeGraph’s Placekey system, a universal location identifier that combines a geospatial encoding system with a unique POI identifier that provides information including the name, latitude and longitude, business details, and industry classification. For privacy reasons, SafeGraph omits sparse data in which fewer than four visitors are recorded in a given week from any home census block group to a POI. Each edge $(i,j)$ in $G(t)$ has a weight $B_{ij}(t)$ that encodes the number of observed movements from $CBG_{i}$ to $POI_{j}$ during week $t$. Equivalently, each graph $G(t)$ can be encoded by a weighted adjacency matrix, $B(t)\in\mathbb{R}^{H\times P}$.
@@
We first discuss the coarse-graining of SafeGraph data to the spatial resolution of census tracts. We define $\mathcal{C}=\{CT_{i}\}_{i=1}^{C}$ to be a set of census tracts of interest, where $C$ is the total number of census tracts ($C=786$ for Harris County) and let $\mathcal{H}_{i}$ and $\mathcal{P}_{i}$ denote, respectively, the census block groups and POIs within $CT_{i}$. Then the combined observed movement from census block groups in $CT_{i}$ to POIs in $CT_{j}$ during week $t$ is given by
@@
The remainder of this study examines time-varying networks encoded by square adjacency matrices $A(t)$ that are size $C\times C$. Finally, for each network, we define $d^{out}_{i}(t)=\sum_{j}A_{ij}(t)$ to be the out-degree—that is, a measure for all movements during week $t$ that leave $CT_{i}$—and $d^{in}_{i}(t)=\sum_{j}A_{ji}(t)$ to be the in-degree—that is, a measure for all movements to POIs within $CT_{j}$ during week $t$. For the mobility network in Figure 1(b), for example, in- and out-degrees for $CT_{1}$ would be $d^{in}_{1}(t)=A_{11}(t)$ and $d^{out}_{1}(t)=A_{11}(t)+A_{12}(t)+A_{13}(t)$.
@@
We next discuss behavioral-stratification of the networks into layers encoding movements to industry categories based on the NAICS heirarchy. Letting $\mathcal{N}$ denote the set of NAICS codes with a fixed number of digits, for each $n\in\mathcal{N}$ we define $\mathcal{P}^{(n)}_{j}$ as the set of POIs within $CT_{j}$ having NAICS code $n$. The network layers’ adjacency matrices are then obtained by $A^{(n)}_{ij}(t)=\sum_{i^{\prime}\in\mathcal{H}_{i},j^{\prime}\in\mathcal{P}^{(n)}_{j}}B_{i^{\prime}j^{\prime}}(t).$ For each network layer $n$ and $CT_{i}$, we define the time-varying in- and out-degrees by $d^{in,(n)}_{i}(t)=\sum_{j}A^{(n)}_{ji}(t)$ and $d^{out,(n)}_{i}(t)=\sum_{j}A^{(n)}_{ij}(t)$, respectively. Finally, note that summing $A^{(n)}_{ij}(t)$ over all possible $n\in\mathcal{N}$ recovers the adjacency matrix $\sum_{n}A^{(n)}(t)=A(t)$ of the original, non-stratified network (i.e., also called the layer-aggregated network). One can similarly obtain the in- and out-degrees for the network encoding all movements by summing over the network layers: $d^{in}_{i}(t)=\sum_{n}d^{in,(n)}_{i}(t)$ and $d^{out}_{i}(t)=\sum_{n}d^{out,(n)}_{i}(t)$.
4.2 Determining threshold for high-movement categories
@@ 4.3 Z-scores quantify movement-change significance
We quantify the storm’s impact on total movements for different network layers by comparing the total movements $m^{(n)}(t)$ during the storm week to a baseline of weekly movement, and we use z-scores to measure deviation from typical behavior. We apply this approach to different network layers to identify which movement categories undergo statistically significant change. After visually inspecting time series encoding weekly movements for several years of data (see Supplementary Figure 2), we select the baseline that consists of the six weeks prior to the storm, $\mathcal{T}_{base}=\{2,3,\dots,7\}$. Letting $\mu$ and $\sigma$ denote the mean and standard deviation of $m(t)$ across $t\in\mathcal{T}_{base}$, we compute the z-score $Z(t)=\frac{m(t)-\mu}{\sigma}$ to study the storm’s impact on all movement categories. Similarly, for each movement category $n$ we let $\mu^{(n)}$ and $\sigma^{(n)}$ denote the mean and standard deviation of $m^{(n)}(t)$ across $t\in\mathcal{T}_{base}$ and compute the z-score
@@
We compare z-scores across movement categories in Section 2.2.
@@ 4.4 Regression analysis relates movements to infrastructure, demographic, and socioeconomic information
In Section 2.5, we study the in- and out-degrees for network layers and investigate how these structural properties relate to infrastructure information (i.e., derived from the POIs) as well social factors including demographic and socioeconomic information. To this end, we gathered data from the U.S. Census Bureau, accessing tables from the 2010 American Community Survey and filtering for year 2019 and census tracts in Harris County, TX. For each census tract, we assembled data for 13 social factors: population (B01003), population density, under 18 (DP05), under 5 (DP05), income (B19013), unemployment (DP03), poverty rate (S0601), non-white percentage (B02001), non-hispanic and non-black percentage (S0601), owner occupied percentage (B25003), renter occupied percentage (B25003), education level (S0601).
To identify which social factors have the strongest correlation with the networks’ in- and out-degrees, we conduct a multivariate linear regression. Noting that some social factors are correlated, provide redundant information, and cause regression instability, we sought to obtain a smaller set of social factors. Specifically, we conducted variance inflation factor (VIF) tests to identify and remove variables having multicollinearity, which can distort regression coefficients and reduce the model’s reliability. For each social factor, $X^{(i)}\in\mathbb{R}^{786},i=1,\dots,13$, we calculated $VIF_{i}=1/(1-R^{2}_{i})$, where
@@
$R^{2}_{i}$ is called the coefficient of determination. Here, each $\hat{X}^{(i)}$ is a predicted value from regression and $\bar{X}^{(i)}$ is the mean across observed values.
In Table 1, we summarize our results for several VIF tests that were were run using the statsmodels module in Python. For Test 1, we used 12 predictors that span three categories: demographic, socioeconomic, and race/ethnicity. In Test 2, we removed a predictor from each category, which were selected based on having a high VIF value in Test 1. For Test 3, we removed two predictors with the highest VIF values from Test 2. Finally, Test 4 showed that the VIF values of all remaining predictors were less than 5 (which is a standard threshold in VIF tests). The final six predictors are: population, population density, income, unemployment %, poverty rate %, and non-white %.
Category Predictor Test 1 Test 2 Test 3 Test 4
Demographic Population 5.318 4.785 3.531 3.375
Population Density 3.308 3.292 2.687 2.672
65 years and over (%) 8.458 5.516 5.444
Under 18 (%) 84.755
Socioeconomic Income 14.488 6.156 5.177 2.675
School Enrollment (%) 63.847 12.666
Unemployment (%) 4.884 4.814 4.535 4.255
Poverty Rate (%) 8.111 7.557 4.021 3.933
Renter Occupied (%) 18.919 8.367
Owner Occupied (%) 28.744
Race/Ethnicity Non-White (%) 6.201 5.609 5.038 4.849
Non-Hispanic and Non-Black (%) 10.992
Table 1: Variance inflation factor (VIF) tests to identify and remove variables having multicollinearity. For Test 1, we used 12 predictors from 3 categories and resulted in most VIF values being greater than 5 (which is a standard threshold). Then we conducted successive tests to remove the highest VIF score from each category (while leaving at least one per category). For the final test (Test 4), each predictor in Test 3 was removed one at a time to identify a set of factors such that no VIF values were greater than 5.
In addition to the social features above, we also include infrastructure features for each census tract in the form of number of POIs. We look at total number of POIs for each census tract as well as a breakdown of number of POIs for the high-movement categories, $\mathcal{N}_{high}$, shown in Figure 2 (c). We highlight that we never include both the total number of POIs and the stratified POIs into categories, since that would introduce collinearity into the model (i.e., the total of POIs equals the summation over POIs in different categories). Each model defines a relationship between either in- or out-degrees and a set of features, $\left\{X^{(i)}\mid i\in S\right\}$, where $S$ is a set of select indexed features (possibly including demographic, socioeconomic, and infrastructure information). The features were standardized using z-score normalization (mean = 0, standard deviation = 1) prior to regression to ensure comparability of coefficients across features. We then fit linear regression models of the form